OpenStudy (anonymous):

Medal! Solve 2x^2 - 8x = -7

OpenStudy (danjs):

DO you recall the quadratic formula thing

OpenStudy (anonymous):

Umm, i recall it i just don't remember it...

OpenStudy (danjs):

for Ax^2 + Bx + C = 0 \[x = \frac{ -B +- \sqrt{B^2 - 4* A *C} }{ 2A }\]

OpenStudy (danjs):

here, you have 2x^2 - 8X + 7 = 0 A = 2 B= -8 C = 7

OpenStudy (anonymous):

Yes. So now just plug it in and solve for that?

OpenStudy (danjs):

right

OpenStudy (anonymous):

x = 8 +-√-8^2 - 4 * 2 * 7 / 2(2)

OpenStudy (anonymous):

well then it doesn't like my square root sign, thats what the question marks are

OpenStudy (danjs):

reload page

OpenStudy (anonymous):

So, then i would start by simplifying the 2(2) and everything inside the square root right?

OpenStudy (danjs):

right so you get \[x = \frac{ 8 +- \sqrt{(-8)^2- 4 * 2 * 7} }{ 2(-8) }\]

OpenStudy (danjs):

\[x = \frac{ 8 +- \sqrt{8} }{ -16 }\]

OpenStudy (anonymous):

Are you sure for the denominator? because A is 2

OpenStudy (danjs):

Yes you are right, my mistake. lol

OpenStudy (danjs):

\[\sqrt{8} = \sqrt{2*4} = 2\sqrt{2}\]

OpenStudy (danjs):

\[\frac{ 8 +-2\sqrt{2} }{ 4 }\]

OpenStudy (danjs):

divide each tem by 2

OpenStudy (danjs):

\[\frac{ 4 +- \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

ok..

OpenStudy (danjs):

\[x = \frac{ 4 + \sqrt{2} }{ 2 } \] or \[x = \frac{ 4 - \sqrt{2} }{ 2 }\]

OpenStudy (danjs):

Thats it, unless you use a calculator to get a decimal number for both of em

OpenStudy (anonymous):

negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2 are my options

OpenStudy (danjs):

Right, if you divide the 2 into both of em you get \[2 +or- \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (danjs):

so the third one

OpenStudy (anonymous):

OK! thanks! and could you explain how you simplified inside the square root for me?

OpenStudy (danjs):

wait, not third

OpenStudy (anonymous):

√(−8)^2−4∗2∗7

OpenStudy (danjs):

the fourth one

OpenStudy (anonymous):

@DanJS thanks for the help but could you tell me how you simplified inside of the square root?

OpenStudy (anonymous):

@DanJS ?

OpenStudy (danjs):

sure

OpenStudy (danjs):

\[\sqrt{(-8)^2 - 4 * 2 *7}\]

OpenStudy (danjs):

if you are squaring a negative number, remember it is (-8)(-8) = +64 not -(8)(8) = -64 The parenthesis go around the negative sign too

OpenStudy (danjs):

\[\sqrt{64 - 56}\]

OpenStudy (danjs):

\[\sqrt{8}\]

OpenStudy (danjs):

\[\sqrt{2 * 4} = \sqrt{ 2}\sqrt{4} = 2*\sqrt{2}\]

OpenStudy (anonymous):

This helps so much thanks!

OpenStudy (danjs):

so square root of 8 = 2*square root of 2

OpenStudy (danjs):

no prob, yw

OpenStudy (anonymous):

@DanJS Ok, so on my next problem, it ends up with a square root of 29 once i simplify it from long form, so how would i do that into something like 4 * 2?

OpenStudy (danjs):

you cant, 29 is a prime number, root 29 is the most you can simplify it

OpenStudy (anonymous):

ok